OpenStudy (anonymous):

solve cosx - sqrt3sinx = 2 a. x = 60° b. x = 300° c. x = 60° , 240° d. x = 0°, 60° , 240°

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \cos x - \sqrt{~3~}\sin x =2 }\) You know that: \(\color{#000000 }{ \displaystyle \cos^2x + \sin^2x=1 }\) \(\color{#000000 }{ \displaystyle \cos^2x =1- \sin^2x }\) \(\color{#000000 }{ \displaystyle \cos x =\sqrt{1- \sin^2x} }\) So, \(\color{#000000 }{ \displaystyle \sqrt{1- \sin^2x}- \sqrt{~3~}\sin x =2 }\) \(\color{#000000 }{ \displaystyle \sqrt{1- \sin^2x} =2+\sqrt{~3~}\sin x }\) \(\color{#000000 }{ \displaystyle (\sqrt{1- \sin^2x})^2 =(2+\sqrt{~3~}\sin x)^2 }\) \(\color{#000000 }{ \displaystyle 1- \sin^2x=(2+\sqrt{~3~}\sin x)^2 }\)

OpenStudy (solomonzelman):

simplify that, and then let sin(x)=a

OpenStudy (solomonzelman):

Then, there is another rule: \(\color{#000000 }{ \displaystyle(a+b)^2=a^2+2ab+b^2 }\) And in this case, \(\color{#000000 }{ \displaystyle (2+\sqrt{~3~}\sin x)^2=2^2+2\cdot 2(\sqrt{~3~}\sin x)+(\sqrt{~3~}\sin x)^2 }\)

OpenStudy (anonymous):

thank you!